# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v2

@@ VI Conclusion
This study tackles the key challenges in infrared small target detection by leveraging background semantics. We introduce DenseSIRST, a benchmark dataset with detailed background annotations, enabling the transition from sparse to dense target detection. Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange. Moreover, we develop BAG-CP, a data augmentation method that selectively pastes targets into sky regions, outperforming the original Copy-Paste method. Extensive experiments validate the effectiveness of BAFE-Net and BAG-CP in improving detection accuracy and reducing false alarms. Our work underscores the importance of explicitly leveraging contextual information in infrared small target detection and paves the way for the development of more robust and context-aware algorithms in this domain.

### v3

@@ VI Conclusion
This study tackles the key challenges in infrared small target detection by leveraging background semantics. We introduce DenseSIRST, a benchmark dataset with detailed background annotations, enabling the transition from sparse to dense target detection. Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange, promoting the exchange of information between the classification and background segmentation branches. Moreover, we develop BAG-CP, a synthesis method for the DenseSIRST dataset, which also serves as a data augmentation technique. BAG-CP selectively pastes targets into sky regions, outperforming the original Copy-Paste method. Extensive experiments validate the effectiveness of BAFE-Net in improving detection accuracy and reducing false alarms. Our work underscores the importance of explicitly leveraging contextual information in infrared small target detection and paves the way for the development of more robust and context-aware algorithms in this domain.